The volume of air at a pressure of 5 bar and at a temperature of 470o C is 0.5 m3. Calculate the mass of the air, and the gas constant. The universal gas constant is 8.315kJ/kg K.


1
Expert's answer
2020-11-24T05:42:26-0500
"PV= nRT"

also "PV= \frac{mass}{m.mass}\cdot RT"

molar mass(air) = 0.029kg/mol

P= pressure= 5 Bar = 500000N/m2

V= volume = 0.5m3

T= Temperature= 470C = 470 + 273= 743K

R= universal gas constant= 8.315kJ/kg

"500000\times0.5 = n\times 8.315\times743"

"n" = "\frac{500000\times0.5}{8.315\times743}" mol

= 40.466mol


"n=\frac{mass}{molar. mass}"

40.466"=\frac{mass}{0.029}"

mass "\approx" 1.174Kg



Gas constant of air

"Gas constant = \frac{universal \hspace{0.05cm}gas\hspace{0.05cm}constant}{m.mass}"


"gas\hspace{0.05cm}constant = \frac{8.315}{0.029}"

"gas constant = 287J/Kg\cdot K"